OpenStudy (anonymous):

Find the volume of a rectangular prism if the length is x, the width is x2, and the height is 5x2 + 4x + 1. Use the formula V = l ⋅ w ⋅ h, where l is length, w is width, and h is height, to find the volume. 5x5 + 4x4 + x3 5x4 + 4x3 + x2 6x5 + 4x4 + x3 6x4 + 4x3 + x2

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

\((x)(x^2)(5x^2+4x+1)\)

Mehek (mehek14):

multiply each number by x

OpenStudy (anonymous):

so c is correct?

Mehek (mehek14):

hold on brb in 5 mins

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

c is correc?

Mehek (mehek14):

no \(x*x^2=x^3\\x^3*5x^2=5x^5\)

Mehek (mehek14):

so the first number in the polynomial is \(5x^5\) in A

OpenStudy (anonymous):

so b is ocrrect / @Mehek14

Mehek (mehek14):

no A